OpenStudy (mpari13):

A motorcycle officer hidden at an intersection observes a car driven by an oblivious driver who ignores a stop sign and continues through the intersection at constant speed. The police officer takes off in pursuit 2.18 s after the car has passed the stop sign. She accelerates at 4.2 m/s2 until her speed is 112 km/h, and then continues at this speed until she catches the car. At that instant, the car is 1.5 km from the intesection. a)how long did it take the officer to cath up to the car? b)how fast was the car traveling (in mi/h) when i worked it out: 9.58seconds and didnt get b)

OpenStudy (amistre64):

hmm, might help to break it into parts. how long does it take the cop to get up to a constant speed, and how far have they traveled? the rest is basic after that.

OpenStudy (amistre64):

looks like we want to convert km to m as well

OpenStudy (amistre64):

a = 4.2 v = 4.2t x = 2.1t^2 112km 1 hr 1000 m ------- * ------ * ------- = 112000m/60s h 60 sec 1 km 11200m/6s = 5600m/3s 5600m s^2 ----------- = 4000/9 seconds so get up to that speed 3s * 4.2 m which is about 7.5 minutes :/ whered i mess up if anything?

OpenStudy (amistre64):

112hm/h is spose to convert to about 31.11 m/s

OpenStudy (amistre64):

lol, there isnt 60 secs in an hour

OpenStudy (amistre64):

thats better .. 280m/9s -------- = 200/27 s, about 7.41 seconds. 4.2 m/s^2 so the cop is at a contant speed at a distance of: x = 2.1(200/27)^2, or about .115 km from the intersection. 2.18 +7.41 ------ 9.59 seconds have passed and they have to catch the car now. 1.500 - 0.115 ------- .385 km to catch up going 112km/h seems fair enough now.

OpenStudy (amistre64):

since d = rt, and we know d and t; r = d/t for the speed of the driver.